# The flexible Gumbel distribution: A new model for inference about the mode

### v2

@@ Abstract
A new unimodal distribution family indexed by the mode and three other parameters is derived from a mixture of a Gumbel distribution for the maximum and a Gumbel distribution for the minimum. Properties of the proposed distribution are explored, including model identifiability and flexibility in capturing heavy-tailed data that exhibit different directions of skewness over a wide range. Both frequentist and Bayesian methods are developed to infer parameters in the new distribution. Simulation studies are conducted to demonstrate satisfactory performance of both methods. By fitting the proposed model to simulated data and data from an application in hydrology, it is shown that the proposed flexible distribution is especially suitable for data containing extreme values in either direction, with the mode being a location parameter of interest. A regression model concerning the mode of a response given covariates based on the proposed unimodal distribution can be easily formulated, which we apply to data from an application in criminology to reveal interesting data features that are obscured by outliers. Computer programs for implementing all considered inference methods in the study are available at https://github.com/rh8liuqy/flexible_Gumbel.
@@ 1 Introduction
Please read the most recent version of the manuscript here. The mean, median, and mode are three most commonly used measure of central tendency of data. When data contain outliers that cause heavy tails or are potentially skewed, the mode is a more sensible representation of the central location of data than the mean or median. The timely review on mode estimation and its application by Chacón [2020] and references therein provide many examples in various fields of research where the mode serves as a more informative representative value of data. Most existing methods developed to draw inference for the mode are semi-/non-parametric in nature, starting from early works on direct estimation in the 1960s [Chernoff, 1964, Dalenius, 1965, Venter, 1967] to more recent works based on kernel density estimation [Chen, 2018] and quantile-based methods [Ota et al., 2019, Zhang et al., 2021]. There are two main reasons contributing to the long-lasting trend of opting to semi-/non-parametric methods for mode estimation, despite the fact that inference procedures proposed along these veins are usually less straightforward to implement (e.g., involving bandwidth selection), and less efficient than their parametric counterparts. First, a parametric model typically imposes stringent constraints on the relationship between the mode and other location parameters that may not be satisfied in a given application. Second, very few existing named distribution families that allow inclusion of both symmetric and asymmetric distributions in the same family can be parameterized so that it is indexed by the mode as the location parameter along with other parameters, such as shape or scale parameters. In this study, we alleviate concerns raised by both reasons that discourage use of parametric methods for mode estimation by formulating a flexible distribution indexed by the (unique) mode and parameters controlling the shape and scale.
@@
When it comes to modeling heavy-tailed data, the Gumbel distribution [Gumbel, 1941] is arguably one of the most widely used models in many disciplines. Indeed, as a case of the generalized extreme value distribution [Jenkinson, 1955], the Gumbel distribution for the maximum (or minimum) is well-suited for modeling extremely large (or small) events that produce heavy-tailed data. For example, it is often used in hydrology to predict extreme rainfall and flood frequency [Loaiciga and Leipnik, 1999, Koutsoyiannis, 2004, Dawley et al., 2019]. In econometrics, the Gumbel distribution plays an important role in modeling extreme movements of stock prices and large changes in interest rates [Bali, 2003, Pratiwi et al., 2019]. The Gumbel distribution is indexed by the mode and a scale parameter, and thus is convenient for mode estimation. However, the Gumbel distribution for the maximum (or minimum) is right-skewed (or left-skewed) with the skewness fixed at around 1.44 (or $-1.44$), and the kurtosis fixed at 5.4 across the entire distribution family. Thus it may be too rigid for scenarios where the direction and extremeness of outliers presented in data are initially unclear, or when the direction and level of skewness are unknown beforehand. Constructions of more flexible distributions that overcome these limitations have been proposed. In particular, Cooray [2010] applied a logarithmic transformation on a random variable following the odd Weibull distribution to obtain the so-called generalized Gumbel distribution that includes the Gumbel distribution as a subfamily. But the mode of the generalized Gumbel distribution is not a location parameter this distribution is indexed by, or an explicit function of other model parameters. Shin et al. [2015] considered mixture distributions with one of the components being the Gumbel distribution and the other component(s) being Gumbel of the same skewness direction or a different distribution, such as the gamma distribution. Besides the same drawback pointed out for the generalized Gumbel distribution, it is difficult to formulate a unimodal distribution following their construction of mixtures, and thus their proposed models are unsuitable when unimodality is a feature required to make inferring the mode meaningful, such as in a regression setting, as in modal regression [Yao et al., 2012, Yao and Li, 2013, Chen, 2018].
@@
With heavy-tailed data in mind and the mode as the location parameter of interest, we construct a new unimodal distribution that does not impose stringent constraints on how the mode relates to other central tendency measures, while allowing a range of kurtosis wide enough to capture heavy tails at either direction, as well as different degrees and directions of skewness. This new distribution, called the flexible Gumbel (FG) distribution, is presented in Section 2, where we study properties of the distribution and discuss identifiability of the model. We present a frequentist method and a Bayesian method for estimating parameters in the FG distribution in Section 3. Finite sample performance of these methods are inspected in simulation study in Section 4, followed by an application of the FG distribution in hydrology in Section 5. Section 6 demonstrates fitting a modal regression model based on the FG distribution to data from a criminology study. Section 7 highlights contributions of the study and outlines future research directions.
@@ 2 The flexible Gumbel distribution
The variance of $Y$ is $V(Y)=\{w\sigma_{1}^{2}+(1-w)\sigma_{2}^{2}\}\pi^{2}/6+w(1-w)(\sigma_{1}+\sigma_{2})^{2}\gamma^{2}$, which does not depend on the mode parameter $\theta$. Obviously, by setting $w=0$ or 1, $\mbox{FG}(\theta,\sigma_{1},\sigma_{2},w)$ reduces to one of the Gumbel components. Unlike a Gumbel distribution that only has one direction of skewness at a fixed level (of ±1.44$\pm$ 1.44)± 1.44 ), an FG distribution can be left-skewed, or right-skewed, or symmetric. More specifically, with the mode fixed at zero when studying the skewness and kurtosis of FG, one can show that the third central moment of $Y$ is given by
@@
where $\bar{w}=1-w$, and $\zeta(3)\approx 1.202$ is Apéry’s constant. Although the direction of skewness is not immediately clear from (4), one may consider a special case with $w=0.5$ where (4) reduces to $(\sigma_{1}-\sigma_{2})\{\gamma\pi^{2}(\sigma_{1}+\sigma_{2})^{2}/8+\zeta(3)(\sigma_{1}^{2}+\sigma_{1}\sigma_{2}+\sigma_{2}^{2})\}$. Now one can see that $\mbox{FG}(\theta,\sigma_{1},\sigma_{2},0.5)$ is symmetric if and only if $\sigma_{1}=\sigma_{2}$, and it is left-skewed (or right-skewed) when $\sigma_{1}$ is less (or greater) than $\sigma_{2}$. The kurtosis of $Y$ can also be derived straightforwardly, with a more lengthy expression than (4) that we omit here, which may not shed much light on its magnitude except for that it varies as the scale parameters and the mixing proportion vary, instead of fixing at 5.4 as for a Gumbel distribution. An R Shiny app depicting the pdf of $\mbox{FG}(\theta,\sigma_{1},\sigma_{2},w)$ with user-specified parameter values is available at https://qingyang.shinyapps.io/gumbel_mixture/, created and maintained by the first author. Along with the density function curve, the Shiny app provides skewness and kurtosis of the depicted FG density. From there one can see that the skewness can be much lower than $-1.44$ or higher than 1.44, and the kurtosis can be much higher than 5.4, suggesting that inference based on FG can be more robust to outliers than when a Gumbel distribution is assumed for data at hand, without imposing stringent assumption on the skewness of the underlying distribution.
@@
The flexibility of a mixture distribution usually comes with concerns relating to identifiability [Teicher, 1961, 1963, Yakowitz and Spragins, 1968]. In particular, there is the notorious issue of label switching when fitting a finite mixture model [Redner and Walker, 1984]. Take the family of two-component normal mixture (NM) distributions as an example, defined by $\{\mbox{NM}(\mu_{1},\sigma_{1},\mu_{2},\sigma_{2},w):\,w\mathcal{N}(\mu_{1},\sigma_{1}^{2})+(1-w)\mathcal{N}(\mu_{1},\sigma_{2}^{2}),\mbox{ for $\sigma_{1},\sigma_{2}>0$ and $w\in[0,1]$}\}$. When fitting a data set assuming a normal mixture distribution, one cannot distinguish between, for instance, $\mbox{NM}(1,2,3,4,0.2)$ and $\mbox{NM}(3,4,1,2,0.8)$, since the likelihood of the data is identical under these two mixture distributions. As another example, for data from a normal distribution, a two-component normal mixture with two identical normal components and an arbitrary mixing proportion $w\in[0,1]$ leads to the same likelihood, and thus $w$ cannot be identified. Teicher [1963] showed that imposing an lexicographical order for the normal components resolves the issue of non-identifiability, which also excludes mixtures with two identical components in the above normal mixture family. Unlike normal mixtures of which all components are in the same family of normal distributions, the FG distribution results from mixing two components from different families, i.e., a Gumbel distribution for the maximum and a Gumbel distribution for the minimum, with weight $w$ on the former component. By construction, FG does not have the label-switching issue. And, according to Teicher [1963, Theorem 1], the so-constructed mixture distribution is always identifiable even when the true distribution is a (one-component) Gumbel distribution.
@@ 3.1 Frequentist inference method
Based on a random sample of size $n$ from the FG distirbution, $\mathbf{y}=\{y_{i}\}_{i=1}^{n}$, maximum likelihood estimators (MLE) of all model parameters in $\mathbf{\Omega}=(\theta,\sigma_{1},\sigma_{2},w)$ can be obtained via the expectation-maximization (EM) algorithm [Dempster et al., 1977]. To apply the EM algorithm, we introduce a latent variable $Z$ that follows Bernoulli($w$) such that the joint likelihood of $(Y,Z)$ is
$f_{Y,Z}(y,z)=\{wf_{1}(y;\theta,\sigma_{1})\}^{z}\{(1-w)f_{2}(y;\theta,\sigma_{2})\}^{1-z},$
where $f_{1}(y;\theta,\sigma_{1})$ is the pdf in (1) evaluated at $y$ with the scale parameter $\sigma=\sigma_{1}$, and $f_{2}(y;\theta,\sigma_{2})$ is the pdf in (2) evaluated at $y$ with the scale parameter $\sigma=\sigma_{2}$. A random sample of size $n$ from Bernoulli($w$), $\mathbf{z}=\{z_{i}\}_{i=1}^{n}$, is viewed as missing data, and $\{(y_{i},z_{i})\}_{i=1}^{n}$ are viewed as the complete data in the EM algorithm. The complete-data log-likelihood is then
@@
In the M-step at the $(t+1)$-th iteration, one maximizes $Q(\boldsymbol{\Omega}|\boldsymbol{\Omega}^{(t)})$ with respect to $\mathbf{\Omega}$ to obtain an updated estimate $\mathbf{\Omega}^{(t+1)}$.
Our experience with the above EM algorithm for fitting the FG distribution suggests that maximizing $Q(\boldsymbol{\Omega}|\boldsymbol{\Omega}^{(t)})$ in (6) can be numerically challenging. We thus exploit the expectation-conditional maximization (ECM) algorithm [Meng and Rubin, 1993], which replaces the M-step with a sequence of simpler conditional maximizations referred to as the CM-step. More specifically, in the CM-step, the updating formula for $w$ is simply $w^{(t+1)}=\sum_{i=1}^{n}T_{i}^{(t)}/n.$ There is no closed-form updating formula for the other three parameters in $\mathbf{\Omega}$, but they can now be easily updated by most well-accepted one-dimensional optimization algorithms, such as the Newton-Raphson method. To ensure convergence at the global maximum, as recommended by Wu [1983], one should implement the ECM algorithm several rounds with different starting values $\mathbf{\Omega}^{(0)}$.
@@ 3.2 Bayesian inference method
$w$ $\sim\text{Beta}(1,1),$
where inv-Gamma refers to the inverse Gamma distribution. We choose the above prior for the scale parameters by following the prior selection for variance parameters suggested in Gelman [2006]
We employ the Metropolis-within-Gibbs sampler [Müller, 1991, 1993] to obtain an estimate of $\mathbf{\Omega}$ from the posterior distribution of $\mathbf{\Omega}$ given observed data $\mathbf{y}$. Similar to the EM/ECM algorithm in Section 3.1, the latent variable $Z$ is also introduced as a device to carry out data augmentation. And the iterative algorithm presented next is based on the following two conditional distributions that can be easily proved,
@@
In Steps 3 and 4, $p(\cdot|\cdot)$ refers to a conditional pdf generically, $\tau_{0}$, $\tau_{1}$, and $\tau_{2}$ are three positive tuning parameters whose values should be chosen so that the acceptance rate at each step is around $23\%$ [Gelman et al., 1997]. To draw samples from the joint posterior distribution, there are numerous ways to design the Markov chain Monte Carlo (MCMC) sampler. Instead of the Metropolis-within-Gibbs sampler we adopt here, one may use other existing MCMC software, such as Stan [Stan Development Team, 2021], JAGS [Plummer et al., 2003], and BUGS [Spiegelhalter et al., 1996, Lunn et al., 2009], two of which we demonstrate in the Appendix. After obtaining enough high quality samples from the joint posterior distribution $p(\theta,\sigma_{1},\sigma_{2},w|\mathbf{y})$, Bayesian inference is straightforward, including point estimation, interval estimation, and uncertainty assessment.

### v3

@@ Abstract
A new unimodal distribution family indexed by the mode and three other parameters is derived from a mixture of a Gumbel distribution for the maximum and a Gumbel distribution for the minimum. Properties of the proposed distribution are explored, including model identifiability and flexibility in capturing heavy-tailed data that exhibit different directions of skewness over a wide range. Both frequentist and Bayesian methods are developed to infer parameters in the new distribution. Simulation studies are conducted to demonstrate satisfactory performance of both methods. By fitting the proposed model to simulated data and data from an application in hydrology, it is shown that the proposed flexible distribution is especially suitable for data containing extreme values in either direction, with the mode being a location parameter of interest. Using the proposed unimodal distribution, one can easily formulate a regression model concerning the mode of a response given covariates. We apply this model to data from an application in criminology to reveal interesting data features that are obscured by outliers. Computer programs for implementing all considered inference methods in the study are available at https://github.com/rh8liuqy/flexible_Gumbel.
@@ 1 Introduction
The mean, median, and mode are the three most commonly used measure of central tendency of data. When data contain outliers that cause heavy tails or are potentially skewed, the mode is a more sensible representation of the central location of data than the mean or median. The timely review on mode estimation and its application by Chacón [2020] and references therein provide many examples in various fields of research where the mode serves as a more informative representative value of data. Most existing methods developed to draw inference for the mode are semi-/non-parametric in nature, starting from early works on direct estimation in the 1960s [Chernoff, 1964, Dalenius, 1965, Venter, 1967] to more recent works based on kernel density estimation [Chen, 2018] and quantile-based methods [Ota et al., 2019, Zhang et al., 2021]. Two main factors contribute to the enduring preference for semi-/non-parametric methods for mode estimation, despite the typically less straightforward implementation and lower efficiency compared to parametric counterparts. First, parametric models often impose strict constraints on the relationship between the mode and other location parameters, which may not hold in certain applications. Second, very few existing named distribution families that allow the inclusion of both symmetric and asymmetric distributions in the same family can be parameterized so that it is indexed by the mode as the location parameter along with other parameters, such as shape or scale parameters. In this study, we alleviate concerns raised by both reasons that discourage the use of parametric methods for mode estimation by formulating a flexible distribution indexed by the (unique) mode and parameters controlling the shape and scale.
@@
When it comes to modeling heavy-tailed data, the Gumbel distribution [Gumbel, 1941] is arguably one of the most widely used models in many disciplines. Indeed, as a case of the generalized extreme value distribution [Jenkinson, 1955], the Gumbel distribution for the maximum (or minimum) is well-suited for modeling extremely large (or small) events that produce heavy-tailed data. For example, it is often used in hydrology to predict extreme rainfall and flood frequency [Loaiciga and Leipnik, 1999, Koutsoyiannis, 2004, Dawley et al., 2019]. In econometrics, the Gumbel distribution plays an important role in modeling extreme movements of stock prices and large changes in interest rates [Bali, 2003, Pratiwi et al., 2019]. The Gumbel distribution is indexed by the mode and a scale parameter, and thus is convenient for mode estimation. However, the Gumbel distribution for the maximum (or minimum) is right-skewed (or left-skewed) with the skewness fixed at around $1.14$ (or $-1.14$), and the kurtosis fixed at 5.4 across the entire distribution family. Thus it may be too rigid for scenarios where the direction and extremeness of outliers presented in data are initially unclear, or when the direction and level of skewness are unknown beforehand. Constructions of more flexible distributions that overcome these limitations have been proposed. In particular, Cooray [2010] applied a logarithmic transformation on a random variable following the odd Weibull distribution to obtain the so-called generalized Gumbel distribution that includes the Gumbel distribution as a subfamily. But the mode of the generalized Gumbel distribution is not indexed by a location parameter, or an explicit function of other model parameters. Shin et al. [2015] considered mixture distributions with one of the components being the Gumbel distribution and the other component(s) being Gumbel of the same skewness direction or a different distribution, such as the gamma distribution. Besides the same drawback pointed out for the generalized Gumbel distribution, it is difficult to formulate a unimodal distribution following their construction of mixtures, and thus their proposed models are unsuitable when unimodality is a feature required to make inferring the mode meaningful, such as in a regression setting, as in modal regression [Yao et al., 2012, Yao and Li, 2013, Chen, 2018].
@@
With heavy-tailed data in mind and the mode as the location parameter of interest, we construct a new unimodal distribution that does not impose stringent constraints on how the mode relates to other central tendency measures, while allowing a range of kurtosis wide enough to capture heavy tails at either direction, as well as different degrees and directions of skewness. This new distribution, called the flexible Gumbel (FG) distribution, is presented in Section 2, where we study properties of the distribution and discuss identifiability of the model. We present a frequentist method and a Bayesian method for estimating parameters in the FG distribution in Section 3. Finite sample performance of these methods is inspected in simulation study in Section 4, followed by an application of the FG distribution in hydrology in Section 5. Section 6 demonstrates fitting a modal regression model based on the FG distribution to data from a criminology study. Section 7 highlights the contributions of our study and outlines future research directions.
@@ 2 The flexible Gumbel distribution
The variance of $Y$ is $V(Y)=\{w\sigma_{1}^{2}+(1-w)\sigma_{2}^{2}\}\pi^{2}/6+w(1-w)(\sigma_{1}+\sigma_{2})^{2}\gamma^{2}$, which does not depend on the mode parameter $\theta$. Obviously, by setting $w=0$ or 1, $\mbox{FG}(\theta,\sigma_{1},\sigma_{2},w)$ reduces to one of the Gumbel components. Unlike a Gumbel distribution that only has one direction of skewness at a fixed level (of ±1.14$\pm$ 1.14)± 1.14 ), an FG distribution can be left-skewed, or right-skewed, or symmetric. More specifically, with the mode fixed at zero when studying the skewness and kurtosis of FG, one can show (as outlined in Appendix A) that the third central moment of $Y$ is given by
@@
where $\bar{w}=1-w$, and $\zeta(3)\approx 1.202$ is Apéry’s constant. Although the direction of skewness is not immediately clear from (4), one may consider a special case with $w=0.5$ where (4) reduces to $(\sigma_{1}-\sigma_{2})\{\gamma\pi^{2}(\sigma_{1}+\sigma_{2})^{2}/8+\zeta(3)(\sigma_{1}^{2}+\sigma_{1}\sigma_{2}+\sigma_{2}^{2})\}$. Now one can see that $\mbox{FG}(\theta,\sigma_{1},\sigma_{2},0.5)$ is symmetric if and only if $\sigma_{1}=\sigma_{2}$, and it is left-skewed (or right-skewed) when $\sigma_{1}$ is less (or greater) than $\sigma_{2}$. The kurtosis of $Y$ can also be derived straightforwardly, with a more lengthy expression than (4) that we omit here, which may not shed much light on its magnitude except that it varies as the scale parameters and the mixing proportion vary, instead of fixing at 5.4 as for a Gumbel distribution. An R Shiny app depicting the pdf of $\mbox{FG}(\theta,\sigma_{1},\sigma_{2},w)$ with user-specified parameter values is available at https://qingyang.shinyapps.io/gumbel_mixture/, created and maintained by the first author. Along with the density function curve, the Shiny app provides skewness and kurtosis of the depicted FG density. From there one can see that the skewness can be much lower than $-1.14$ or higher than $1.14$, and the kurtosis can be much higher than 5.4, suggesting that inference based on FG can be more robust to outliers than when a Gumbel distribution is assumed for data at hand, without imposing stringent assumption on the skewness of the underlying distribution.
@@
The flexibility of a mixture distribution usually comes with concerns relating to identifiability [Teicher, 1961, 1963, Yakowitz and Spragins, 1968]. In particular, there is the notorious issue of label switching when fitting a finite mixture model [Redner and Walker, 1984]. Take the family of two-component normal mixture (NM) distributions as an example, defined by $\{\mbox{NM}(\mu_{1},\sigma_{1},\mu_{2},\sigma_{2},w):\,w\mathcal{N}(\mu_{1},\sigma_{1}^{2})+(1-w)\mathcal{N}(\mu_{1},\sigma_{2}^{2}),\mbox{ for $\sigma_{1},\sigma_{2}>0$ and $w\in[0,1]$}\}$. When fitting a data set assuming a normal mixture distribution, one cannot distinguish between, for instance, $\mbox{NM}(1,2,3,4,0.2)$ and $\mbox{NM}(3,4,1,2,0.8)$, since the likelihood of the data is identical under these two mixture distributions. As another example, for data from a normal distribution, a two-component normal mixture with two identical normal components and an arbitrary mixing proportion $w\in[0,1]$ leads to the same likelihood, and thus $w$ cannot be identified. Teicher [1963] showed that imposing a lexicographical order for the normal components resolves the issue of non-identifiability, which also excludes mixtures with two identical components in the above normal mixture family. Unlike normal mixtures of which all components are in the same family of normal distributions, the FG distribution results from mixing two components from different families, i.e., a Gumbel distribution for the maximum and a Gumbel distribution for the minimum, with weight $w$ on the former component. By construction, FG does not have the label-switching issue. And, we show in Appendix B by invoking Theorem 1 in Teicher [1963] that the so-constructed mixture distribution is always identifiable even when the true distribution is a (one-component) Gumbel distribution.
@@ 3.1 Frequentist inference method
Based on a random sample of size $n$ from the FG distribution, $\mathbf{y}=\{y_{i}\}_{i=1}^{n}$, maximum likelihood estimators (MLE) of all model parameters in $\mathbf{\Omega}=(\theta,\sigma_{1},\sigma_{2},w)$ can be obtained via the expectation-maximization (EM) algorithm [Dempster et al., 1977]. To apply the EM algorithm, we introduce a latent variable $Z$ that follows Bernoulli($w$) such that the joint likelihood of $(Y,Z)$ is
$f_{Y,Z}(y,z)=\{wf_{1}(y;\theta,\sigma_{1})\}^{z}\{(1-w)f_{2}(y;\theta,\sigma_{2})\}^{1-z},$ (5)
where $f_{1}(y;\theta,\sigma_{1})$ is the pdf in (1) evaluated at $y$ with the scale parameter $\sigma=\sigma_{1}$, and $f_{2}(y;\theta,\sigma_{2})$ is the pdf in (2) evaluated at $y$ with the scale parameter $\sigma=\sigma_{2}$. A random sample of size $n$ from Bernoulli($w$), $\mathbf{z}=\{z_{i}\}_{i=1}^{n}$, is viewed as missing data, and $\{(y_{i},z_{i})\}_{i=1}^{n}$ are viewed as the complete data in the EM algorithm. It can be shown [Boos and Stefanski, 2013, Section 2.6.3a] that integrating $z$ out from (5) indeed gives the density of $Y$ in (3). The log-likelihood based on the density in (3) is usually not well-behaved as an objective function to be maximized with respect to $\mathbf{\Omega}$. By considering the complete-data log-likelihood based on (5), one can often maximize an objective function that is better-behaved as we demonstrate next. More specifically, the complete-data log-likelihood is
@@
In the M-step at the $(t+1)$-th iteration, one maximizes $Q(\boldsymbol{\Omega}|\boldsymbol{\Omega}^{(t)})$ with respect to $\mathbf{\Omega}$ to obtain an updated estimate $\mathbf{\Omega}^{(t+1)}=(\theta^{(t+1)},\sigma_{1}^{(t+1)},\sigma_{2}^{(t+1)},w^{(t+1)})$, in which $w^{(t+1)}=\sum_{i=1}^{n}T_{i}^{(t)}/n$, and the other three updated estimates in $\mathbf{\Omega}^{(t+1)}$ are obtained numerically.
The EM algorithm avoids directly maximizing the log-likelihood based on (3) by (iteratively) maximizing the better-behaved $Q(\mathbf{\Omega}|\mathbf{\Omega}^{(t)})$ in (7). To further improve the numerical efficiency, we exploit the expectation-conditional maximization (ECM) algorithm [Meng and Rubin, 1993], which replaces the M-step with a sequence of simpler conditional maximizations referred to as the CM-step. Essentially, within each $M$-step, we update $w$ via $w^{(t+1)}=\sum_{i=1}^{n}T_{i}^{(t)}/n$, then we update $\theta$ using $w^{(t+1)}$ along with $(\sigma_{1}^{(t)},\sigma_{2}^{(t)})$, followed by updating $\sigma_{1}$ using $w^{(t+1)}$, the recently updated $\theta$, and $\sigma_{2}^{(t)}$; lastly, we update $\sigma_{2}$ using $w^{(t+1)}$ and the recently updated $\theta$ and $\sigma_{1}$. There is no closed-form updating formula for the latter three updates, but each of them can now be easily updated by most well-accepted one-dimensional optimization algorithms, such as the Newton-Raphson method. To ensure convergence at the global maximum, as recommended by Wu [1983], one should implement the ECM algorithm several rounds with different starting values $\mathbf{\Omega}^{(0)}$.
@@ 3.2 Bayesian inference method
[…]
where inv-Gamma refers to the inverse Gamma distribution. We choose the above prior for the scale parameters by following the prior selection for variance parameters suggested in Gelman [2006].
We employ the Metropolis-within-Gibbs sampler [Müller, 1991, 1993] to obtain an estimate of $\mathbf{\Omega}$ from the posterior distribution of $\mathbf{\Omega}$ given observed data $\mathbf{y}$. Similar to the EM/ECM algorithm in Section 3.1, the latent variable $Z$ is also introduced as a device to carry out data augmentation. The iterative algorithm presented next is based on the following two conditional distributions that can be easily proved,
@@
In Steps 3 and 4, $p(\cdot|\cdot)$ refers to a conditional pdf generically, $\tau_{0}$, $\tau_{1}$, and $\tau_{2}$ are three positive tuning parameters whose values should be chosen so that the acceptance rate at each step is around $23\%$ [Gelman et al., 1997]. To draw samples from the joint posterior distribution, there are numerous ways to design the Markov chain Monte Carlo (MCMC) sampler. Instead of the Metropolis-within-Gibbs sampler we adopt here, one may use other existing MCMC software, such as Stan [Stan Development Team, 2021], JAGS [Plummer et al., 2003], and BUGS [Spiegelhalter et al., 1996, Lunn et al., 2009], the former two of which are demonstrated in the GitHub repository (https://github.com/rh8liuqy/flexible_Gumbel). After obtaining enough high-quality samples from the joint posterior distribution $p(\theta,\sigma_{1},\sigma_{2},w|\mathbf{y})$, Bayesian inference is straightforward, including point estimation, interval estimation, and uncertainty assessment.